OpenStudy (anonymous):

factor t^3+3t^2-12t+4

OpenStudy (michele_laino):

hint: we can factor out t^2 between the first and second term, so we get: \[\begin{gathered} {t^3} + 3{t^2} - 12t + 4 = \hfill \\ \hfill \\ = {t^2}\left( {t + 3} \right) - 12t + 4 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

no, I'm sorry, we have to apply another procedure

OpenStudy (anonymous):

okay which one? I was trying to factor by grouping too, but it didn't work for me either

OpenStudy (michele_laino):

we can see that such polynomial is divisible by 2, since \[P\left( 2 \right) = 0\] where \[P\left( t \right) = {t^3} + 3{t^2} - 12t + 4\]

OpenStudy (michele_laino):

so I apply the Ruffini's method: 

OpenStudy (michele_laino):

oops...

OpenStudy (anonymous):

wait where did the 2 come from? how did you get it? how do you know that p(2)=0?

OpenStudy (michele_laino):

if I replace t=2 into the polynomial you provided, I get: \[\begin{gathered} P\left( 2 \right) = {2^3} + 3 \cdot {2^2} - 12 \cdot 2 + 4 = \hfill \\ \hfill \\ = 8 + 12 - 24 + 4 = 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, according to the Ruffini's theorem, such polynomial, is divisible by 2, and we can write this: \[\Large {t^3} + 3{t^2} - 12t + 4 = \left( {t - 2} \right)\left( {{t^2} + 5t - 2} \right)\]

OpenStudy (anonymous):

oooh thank you! I will learn about Ruffini's theorem

OpenStudy (michele_laino):

next we have to factor this trinomial: \[{{t^2} + 5t - 2}\]

OpenStudy (michele_laino):

such trinomial, can be factored using this formula: \[\Large {t^2} + 5t - 2 = \left( {t - {t_1}} \right)\left( {t - {t_2}} \right)\] where \t_1,\;t_2\) are the zeros of such trinomal: \[\Large {t^2} + 5t - 2 = 0\]

OpenStudy (michele_laino):

oops.. I meant \(t_1,\;t_2\)

OpenStudy (michele_laino):

please try to solve this quadratic equation: \[\Large {t^2} + 5t - 2 = 0\]

OpenStudy (anonymous):

that's okay, it is factored enough for the limit i solved. Thank you

OpenStudy (michele_laino):

:)